OpenStudy (krazzikitten69):

Help PLEASE? What is the solution of the system? 1. 4x-y= -3 6x - y= -8 -(5,5.5) -(-2.5,-7) -(-2.5,2.2) -(2.5,-7) 2. 5x + 7y= 32 8x + 6y= 46 -(8,5) -(1,5) -(7,0) -(5,1)

OpenStudy (anonymous):

what class is this?

OpenStudy (krazzikitten69):

Algebra 1

OpenStudy (anonymous):

ok im in algebra 2 gimme a sec

OpenStudy (krazzikitten69):

Alright

OpenStudy (anonymous):

now are your answers being multiplied my (-) or is that just hashmarks showing each individual answer?

OpenStudy (krazzikitten69):

Just showing the answers

OpenStudy (anonymous):

Oh ok i was gonna say, this is complicated algebra 1 stuff. ok im gonna show you how to do the first then let you try the second. does that sound good?

OpenStudy (krazzikitten69):

Yes

OpenStudy (anonymous):

so here is what you do to find your answer. first take your first equation: 4x-y= -3 then pick any of the answers, so lets go with A: (5,5.5) plug-in your co-ordinates. 4(5)-(5.5)=-3 multiply: 20-5.5=-3 since 14.5 does not equal -3, you go with the next one. think you got it?

OpenStudy (anonymous):

Then if a answer works, make sure to test it on both equations before choosing it

OpenStudy (anonymous):

so for the first one your answer would be B

OpenStudy (anonymous):

@krazzikitten69 are you there?